### v1

# The Conjugate Function Method for Surfaces with Elaborate Topological Types

Abstract
The conjugate function method is an algorithm for numerical computation of conformal mappings for simply and multiply connected domains on surfaces. In this paper the conjugate function method is generalized and refined to achieve the same level of accuracy on simply and multiply connected planar domains and Riemann surfaces. The main challenge addressed here is the accurate and efficient construction of the conjugate problem for multiply connected domains. The method relies on high-order finite element methods which allow for highly accurate computations of mappings on surfaces, including domains of complex boundary geometry containing strong singularities and cusps. The efficacy of the proposed method is illustrated via an extensive set of numerical experiments with error estimates.
keywords:
Laplace–Beltrami, numerical conformal mappings, conformal modulus, Riemann surfaces
{AMS}
30Cxx, 30Fxx, 65E10, 65N30
1 Introduction
Conformal geometry has many applications such as engineering (e.g. [5]) and mathematical geodesy (see [27]). Traditionally, these applications have been mostly restricted on conformal mappings between planar domains. Even in this setting finding mappings in the case of multiply connected domains has remained challenging. M.M.S. Nasser has used the boundary integral equation (BIE) with impressive results [19, 21]. Moreover, in a very recent work, Nasser discusses how to implement the generalized Koebe’s iterative method efficiently [15].
Koebe’s iterative method can also be implemented for surfaces. Gu, Luo, and Yau give an excellent overview of the state of the art in [6]. Due to both theoretical and implementation complexity, there are relatively few publications addressing specifics. See, for instance, [16, 17, 18, 25, 28]. Even though a properly implemented Koebe’s iteration converges reasonably fast (at least in the planar case), it is very difficult to provide any reliable error estimates or convergence rates.
Our method of choice for conformal mappings, the conjugate function method, was first introduced in [8] for simply and doubly connected planar domains. We use the $hp$-adaptive finite element method ($hp$-FEM) introduced in [11] for computing potential functions and conformal moduli, but other methods can be used as well, e.g., [13] the stochastic version of the algorithm was given for circular arc domains in the plane. One advantage of $hp$-FEM is that in error analysis than not only follows from the adoption of the finite element method for implementation but is also inherent in the form of the reciprocal identity relating the conformal capacities of the primary and conjugate problems. In this work this concept is made precise for surfaces in the specific setting of Q-type canonical domains.
On method development the main contribution of this work is the new, direct algorithm for the construction of the conjugate problem for multiply connected problems. Given the finite element discretization of the primary problem, the problem of finding the right boundary conditions for the conjugate problem, that is, the necessary Dirichlet-Neumann swap of the boundary conditions, can be reduced to a quadratic minimization problem that can be solved with a linear system of equations whose dimension is the number of the holes. Not only is the approach much faster than the previous ones, but the accuracy of the solution of the conjugate problem is brought to the same level as that of the primary problem. The same algorithm applies to problems on surfaces as well without any modifications. The conjugate function method is now comparable to the state-of-the-art methods with the additional benefits of the high-order finite element frameworks such as advanced error estimation and adaptivity.
1.1 Illustrative Example: 3D Face Recognition
Face recognition has many applications, especially in biometrics. Since a standard 2D image recognition is fragile due to many factors, 3D methods have been proposed as a more reliable alternative. Each 3D surface with suitable topology can be mapped conformally to a 2D domain and thus, the image recognition becomes a conformal map comparison problem. See [26].
In Figure 1 one example of how such a process might work is illustrated. First, four points on the outer boundary are fixed. In our formulation the canonical domain is a quadrilateral with horizontal slits if holes are present. Second, the Laplace-Beltrami equation is solved twice with different boundary conditions. As mentioned above, this process is only marginally slower if multiple holes are present. In the example both eyes and the mouth are modelled as holes.
Figure 1: Map of a face. Left: The map on the parameter space. Right: Checkerboard illustration of the map on the surface. Original data source: [24].
1.2 Organization
The rest of the paper is organized as follows: After the introduction, preliminaries including the finite element method used and the planar conjugate function method are covered.
2 Preliminaries
In section some basic concepts and definitions used throughout this paper are introduced.
2.1 Riemann surfaces
Recall that a Hausdorff space $X$ is locally Euclidean of dimension $n\geq 1$ if every point $p$ in $X$ has a neighbourhood $U$ such that there is a homeomorphism $\varphi$ from $U$ onto an open subset of $\mathbb{R}^{n}$. We call the pair $(U,\ \varphi:U\to\mathbb{R}^{n})$ a chart, $U$ a coordinate neighbourhood or a coordinate open set, and $\phi$ a coordinate mapping or a coordinate system on $U$. We say that a chart $(U,\varphi)$ is a centered at $p\in U$ if $\varphi(p)=0$. Let $X$ be a two-dimensional manifold, which we also consider as a one-dimensional complex manifold. A complex chart on $X$ is a homeomorphisms $\varphi:U\to V$ from an open subset $U\subset X$ to an open subset $V\subset\mathbb{C}$. Two complex charts $\varphi_{j}:U_{j}\to V_{j}$ (for $j=1,2$) are said to be holomorphically compatible if the mapping
$\varphi_{2}\circ\varphi_{1}^{-1}:\varphi_{1}(U_{1}\cap U_{2})\to\varphi_{2}(U_{1}\cap U_{2})$
is conformal (biholomorphic). The collection of compatible charts covering the whole space $X$ is called an atlas of the manifold. We call manifolds with an atlas of holomorphically compatible charts a Riemann surface.
The genus of a surface is defined as the largest number of non-intersecting simple closed curves that can be drawn on the surfaces without separating it. Intuitively, this corresponds to the number of holes on the surface. Note that in this paper we consider both closed surfaces and surfaces with multiple boundary components.
For more information of Riemann surfaces, we refer to [4].
2.2 High-Order Finite Element Method
High-order finite element methods have the capability for exponential convergence provided the discretization is constructed properly in both domain and in local polynomial order ($hp$-version). In many cases one has to work with cases where exact parameterization of the boundaries on the parameter space is not known. Then only algebraic convergence rate is to be expected.
Let us consider the Dirichlet-Neumann problem on some polygonal domain and denote its weak solution $u_{0}$. The optimal rate of convergence was first proved by Babuška and Guo [2, 3]. For rigorous treatment of the theory involved see Schwab [23] and references therein.
Theorem 2.1.
Let $\Omega\subset\mathbb{R}^{2}$ be a polygon, $v$ the FEM-solution of the Dirichlet-Neumann problem on some polygonal domain, and let the weak solution $u_{0}$ be in a suitable countably normed space where the derivatives of arbitrarily high order are controlled. Then
$\inf_{v}\|u_{0}-v\|_{H^{1}(\Omega)}\leq C\,\exp(-b\sqrt[3]{N}),$
where $C$ and $b$ are independent of $N$, the number of degrees of freedom. Here $v$ is computed on a proper geometric mesh, where the order of an individual element is set to be its element graph distance to the nearest singularity. (The result also holds for meshes with constant polynomial degree.)
2.2.1 Error Estimation
The solution method specific a posteriori error estimation method applied here is the so-called auxiliary subspace error estimation. The estimate measures the projection of the residual to the auxiliary space, that is, to a set of finite element degrees of freedom that extend the approximation space in a natural way. One example of such extension is to consider all edge modes of degree $p+1$ and bubble modes of $p+1$ and $p+2$, where $p$ is the (constant) polynomial degree used in the discretization. For details, see [7].
First, $p$-robustness has never been rigorously shown for this class of error estimators. Yet, there exists compelling numerical evidence that the method is indeed $p$-robust. Second, by construction, these estimators are optimistic. If properly constructed, they cannot overestimate the error.
2.2.2 Special Types of Boundaries
The computational domains can be defined in different ways, either exactly in some parameterized form or using some given discretization. Our approach for geometric grading of the meshes with known locations of singularities is based on rule based algorithms [12].
One non-standard boundary type in the context of finite element method considered here is the slit or cut. We define a cut to be a set of element nodes and edges that form a loop with zero area. That is, the data structures support having two edges on top of each other, both with their own kinematic constraints.
3 The Conjugate Function Method On Planar Domains and Riemann Surfaces
Let $\Omega\subset\mathcal{S}$ is a simply connected domain on a Riemann surface $\mathcal{S}$ so that its boundary of $\Omega$ is a Jordan curve. We call $\Omega$ together with four positively oriented points $z_{1},z_{2},z_{3},z_{4}\in\partial\Omega$ a (generalized) quadrilateral and denote it by $Q=(\Omega;z_{1},z_{2},z_{3},z_{4})$. The boundary segments connecting the pairs of points $(z_{j},z_{j+1})$ for $j=1,2,3$, and $(z_{4},z_{1})$ for $j=4$, respectively, are denoted by $\gamma_{j}$.
It is well-known (see e.g. [1]) that there exists a unique number $h>0$ called the conformal modulus of $Q$, such that there exists a conformal mapping of the rectangle $R_{h}=[0,1]\times[0,h]\subset\mathbb{C}$ onto $\Omega$, with boundary points $z_{1},z_{2},z_{3},z_{4}$ corresponding to the images of the points $0,1,1+ih,ih$, respectively. The conformal modulus determines the conformal equivalence class of $\Omega$ in the sense that there exists a conformal mapping between quadrilaterals (with boundary point correspondence) if and only if they have the same modulus. In the following the conformal modulus of a quadrilateral $Q$ is denoted by $M(Q)$.
In this study, the conformal modulus of a quadrilateral is computed via its connection to the Laplacian. Recall that there exists a (unique) harmonic solution $u$ to the following Dirichlet-Neumann mixed boundary value problem:
(1) $\left\{\begin{array}[]{rcl}\Delta_{S}u(z)=0&\text{for}&z\in\Omega,\\ u(z)=0&\text{for}&z\in\gamma_{2},\\ u(z)=1&\text{for}&z\in\gamma_{4},\\ \partial u(z)/\partial n=0&\text{for}&z\in\gamma_{1}\cup\gamma_{3},\end{array}\right$.
where $n$ is the unit exterior boundary normal and $\Delta_{S}$ is the Laplace-Beltrami operator. Then the conformal modulus is connected to the above boundary value problem by the identity (see e.g. Ahlfors [1, Theorem 4.5] and Papamichael and Stylianopoulos [22, Theorem 2.3.3]):
(2) $M(Q)=\iint_{\Omega}|\nabla u|^{2}\,dx\,dy.$
For a quadrilateral $Q=(\Omega;z_{1},z_{2},z_{3},z_{4})$ we call $\widetilde{Q}=(\Omega;z_{2},z_{3},z_{4},z_{1})$ its conjugate quadrilateral and the corresponding problem (1) for quadrilateral $\widetilde{Q}$ the conjugate Dirichlet–Neumann problem. It is well-known that if $M(Q)=h>0$, then $M(\widetilde{Q})=1/h$, which leads to the following very useful reciprocal identity: For all quadrilaterals $Q$, we have
(3) $M(Q)M(\widetilde{Q})=1.$
Furthermore, we may observe that canonical conformal mapping of a quadrilateral $Q=(\Omega;z_{1},z_{2},z_{3},z_{4})$ onto the rectangle $R_{h}$ with vertices at $1+ih$, ihih, 0, and $1$, can be obtained by solving the corresponding Dirichlet–Neumann problem $u$ and its conjugate problem.
Lemma 3.1.
Let $Q$ be a quadrilateral with modulus $h$, and suppose $u$ solves the Dirichlet–Neumann problem (1). If $v$ is a harmonic function conjugate to $u$, satisfying $v(\operatorname{Re}z_{3},\operatorname{Im}z_{3})=0$, and $\tilde{u}$ represents the harmonic solution for the conjugate quadrilateral $\widetilde{Q}$, then $v=h\tilde{u}$.
3.1 Q-type canonical domains
Let $h>0$, and let $\big{(}(\zeta_{1},d_{1}),\ldots,(\eta_{N},d_{N})\big{)}$ be complex numbers such that $\operatorname{Re}\zeta_{i}\in(0,1)$ and $\operatorname{Im}\zeta_{i}\in(0,h)$ for all $i=1,\ldots,N$, where $N\geq 1$. For a multiply connected plane case, we will consider canonical domains of the type
(4) $\big{(}(0,1)\times(0,h)\big{)}\setminus\bigcup_{j=1}^{N}[\zeta_{j},\zeta_{j}+d_{j}],$
i.e., a rectangle $(1,0)\times(0,h)$ with $N$ horizontal slits removed. Canonical domains of this type are called quadrilateral-like (or Q-type) in [9].
3.2 Variational Formulation for Finite Elements
Our task is to define the Laplacian on some surface $S$, that is, we want to define the operator $\Delta_{S}$ in the form which is suitable for finite element implementation. In our setting, the surface is always assumed to be given in some parameterised form. Let $\mathbf{x}_{S}:\Gamma\to S$ be a parameterisation of a surface $S$. The goal is to treat $\Gamma\subset\mathbb{R}^{2}$ as the reference domain on which the finite elements are defined. Let $J_{\mathbf{x}}$ be the Jacobian of the mapping, and hence $G_{S}=J_{\mathbf{x}}^{T}J_{\mathbf{x}}$ is the first fundamental form.
The tangential gradient of some function $v:S\to\mathbb{R}$ is
(5) $(\nabla_{S}v)\circ\mathbf{x}_{S}:=J_{\mathbf{x}}G_{S}^{-1}\nabla(v\circ\mathbf{x}_{S}),$
and immediately, using the same notation, the $\Delta_{S}$ can be written as $\Delta_{S}:=\nabla_{S}\cdot\nabla_{S},$ and thus, the variational formulation $\int_{S}\nabla_{S}\psi\cdot\nabla_{S}v\,dx,\ \mbox{ for all }v\in H^{1}(S)$ on an image $K$ of a given element $T$ in discretisation of $\Gamma$ becomes
(6) $\int_{K}\nabla_{K}\psi\cdot\nabla_{K}v\,dx$ $=\int_{T}\nabla(\psi\circ\mathbf{x}_{K})^{T}G_{K}^{-T}J_{\mathbf{x}}^{T}J_{\mathbf{x}}G_{K}^{-1}\nabla(v\circ\mathbf{x}_{K})\sqrt{\operatorname{det}(G_{K})}d\tau$
(7) $=\int_{T}\nabla(\psi\circ\mathbf{x}_{K})^{T}G_{K}^{-T}G_{S}G_{K}^{-1}\nabla(v\circ\mathbf{x}_{K})\sqrt{\operatorname{det}(G_{K})}d\tau.$
The integrals are evaluated on standard 2D mapped Gaussian quadratures. In fact, it is the first equality (6) that is compatible with our implementation of the method. Since the problem has been transformed to a standard 2D planar problem with variable coefficients, there is no need for additional arguments on the convergence of the method.
4 Fast Construction of the Conjugate Problem
The starting point is the quadrilateral $Q$. In the multiply connected case there may be a finite number of holes each with its own boundary $\partial E_{i}$. Initially, each interior boundary will have a zero Neumann boundary condition: $\partial u/\partial n=0$ on every $\partial E_{i}$. The task is to construct the conjugate problem where each of the $\partial E_{i}$ is set to some constant potential $v_{i}$, in other words, a corresponding Dirichlet boundary condition is defined, v=vi, $v_{i},\text{ on }\partial E_{i}$,\ $\forall$ i.
The overall solution process can be compressed to three steps: 1. Solve the primary problem, 2. Construct the conjugate problem, 3. Solve the conjugate problem. That is, considering the primary problem
[…]
These unknown potentials $v_{i}$ can be found using the reciprocal identity discussed in the next subsection. In our previous work, this task has been completed using optimization, which has not only been computationally expensive but has also led to unavoidable loss of accuracy when quadratic objective functions have been used. Remarkably, using the matrix representation of the discretization of the problem and the reciprocal identity, one can derive a two-step method for finding the minimizing Dirichlet boundary conditions for the conjugate problem.
4.1 Reciprocal Identity for Multiply Connected Domains
We note that when the boundary conditions given by (8) are applied to a Q-type canonical domain given by (4), then we obtain the following natural generalization of the modulus of quadrilateral.
Let $\Omega$ be a domain of the type (4) on an oriented surface $S$. Denote by $Q_{*}=(\Omega;z_{1},z_{2},z_{3},z_{4})$ the Q-type domain with points $z_{1},z_{2},z_{3},z_{4}$ chosen in a positive order from the same boundary component. Then the conformal modulus of $Q_{*}$ and its modulus and $\tilde{Q}_{*}=(\Omega;z_{2},z_{3},z_{4},z_{1})$, the conjugate modulus, are defined by
(9) $M(Q_{*})=\iint_{\Omega}|\nabla u|^{2}\,dx\,dy,\quad M(\tilde{Q}_{*})=\iint_{\Omega}|\nabla v|^{2}\,dx\,dy,$
where $u,v$, respectively, are harmonic solutions to the boundary value problems given by (8).
These definitions are natural generalizations of the conformal modulus of a quadrilateral for multiply connected domains, and the classical definition discussed in Section 3 can be understood as their special cases. Furthermore, they admit the following very useful generalization of the reciprocal identity (3):
Proposition 4.1.
Let $\Omega$ be a domain on an oriented surface $S$ that can mapped conformally onto a Q-type canonical domain $Q_{*}=(\Omega;z_{1},z_{2},z_{3},z_{4})$ defined by (4), where the points $z_{1},z_{2},z_{3},z_{4}$ are chosen from the same boundary component of $\Omega$. Then the generalized conformal moduli of $\Omega$ satisfy the reciprocal identity:
(10) $M(Q_{*})M(\tilde{Q}_{*})=1.$
Proof 4.2.
First recall that the Dirichlet energy integral
(11) $\iint_{\Omega}|\nabla u|^{2}\,dx\,dy$
in (2) is a well-known conformal invariant in the plane, and this property extends also to conformal mappings between Riemann surfaces (see, e.g., [14, Proposition 2.1.1, p. 46]). Because the boundary data in (8) is preserved under conformal mappings, it is sufficient to show that this identity is true for the canonical domains given by (4).
But in this case, the solution of the boundary value problem (8) is given by the harmonic functions $u(x,y)=x$ and $v(x,y)=y/h$. Therefore, the desired identity follows immediately.
4.2 Linear Algebra Based Construction of the Conjugate Problem
Let us divide the degrees of freedom, internal and boundary, into sets. The sets of a discretized quadrilateral are denoted $B,D_{0},D_{1},N^{0}$, and $N^{1}$, for internal, Dirichlet $u=0$, Dirichlet $u=1$, Neumann with Dirichlet $u=0$ in the primary problem, and Neumann with Dirichlet $u=1$ in the primary problem, degrees of freedom, respectively. The degrees of freedom on the boundaries of the holes $E_{i}$ are all Dirichlet, constant over the boundary and initially unknown $\delta_{E_{i}}$.
The discretized system is
$A=\begin{pmatrix}A_{BB}&A_{BN^{1}}&A_{BN^{0}}&A_{BD_{1}}&A_{BD_{0}}&A_{BE_{1}}&\dots&A_{BE_{n}}\\ A_{N^{1}B}&A_{N^{1}N^{1}}&A_{N^{1}N^{0}}&A_{N^{1}D_{1}}&A_{N^{1}D_{0}}&A_{N^{1}E_{1}}&&A_{N^{1}E_{n}}\\ A_{N^{0}B}&A_{N^{0}N^{1}}&A_{N^{0}N^{0}}&A_{N^{0}D_{1}}&A_{N^{0}D_{0}}&A_{N^{0}E_{1}}&&A_{N^{0}E_{n}}\\ A_{D_{1}B}&A_{D_{1}N^{1}}&A_{D_{1}N^{0}}&A_{D_{1}D_{1}}&A_{D_{1}D_{0}}&A_{D_{1}E_{1}}&&A_{D_{1}E_{n}}\\ A_{D_{0}B}&A_{D_{0}N^{1}}&A_{D_{0}N^{0}}&A_{D_{0}D_{1}}&A_{D_{0}D_{0}}&A_{D_{0}E_{1}}&&A_{D_{0}E_{n}}\\ A_{E_{1}B}&A_{E_{1}N^{1}}&A_{E_{1}N^{0}}&A_{E_{1}D_{1}}&A_{E_{1}D_{0}}&A_{E_{1}E_{1}}&&A_{E_{1}E_{n}}\\ \vdots&&&&&&\ddots\\ A_{E_{n}B}&A_{E_{n}N^{1}}&A_{E_{n}N^{0}}&A_{E_{n}D_{1}}&A_{E_{n}D_{0}}&A_{E_{n}E_{1}}&\dots&A_{E_{n}E_{n}}\\ \end{pmatrix}$.
The solution vector $x$ (or the potential $v$) has a similar structure
$x=\begin{pmatrix}x_{B}&x_{N^{1}}&x_{N^{0}}&x_{D_{1}}&x_{D_{0}}&x_{E_{1}}&\dots&x_{E_{n}}\end{pmatrix}^{T}.$
In order to derive the quadratic form we first divide the degrees of freedom into free, that is, zero Neumann, and fixed, where the internal and Neumann degrees of freedom are free, and all Dirichlet ones are fixed. These sets are denoted by $I$ and $D$, respectively. The matrix $A$ has a simple $2\times 2$ form $A=\left(\begin{smallmatrix}A_{II}&A_{ID}\\ A_{DI}&A_{DD}\\ \end{smallmatrix}\right)$, and $x=\left(\begin{smallmatrix}x_{I}&x_{D}\end{smallmatrix}\right)^{T}$. Using the block form one can solve (formally) the free degrees of freedom
$A_{II}x_{I}=-A_{ID}x_{D},\quad x_{I}=-A_{II}^{-1}A_{ID}x_{D}.$
Using this, one can then evaluate the quadratic form
(12) $Q_{F}=\begin{pmatrix}x_{I}\\ x_{D}\\ \end{pmatrix}^{T}\begin{pmatrix}A_{II}&A_{ID}\\ A_{DI}&A_{DD}\\ \end{pmatrix}\begin{pmatrix}x_{I}\\ x_{D}\\ \end{pmatrix}=x_{D}^{T}A_{DD}x_{D}-x_{D}^{T}A_{DI}A_{II}^{-1}A_{ID}x_{D}.$
Let us next denote the vector of unknown parameter values $x^{\prime}_{D}=\left(\begin{smallmatrix}\delta_{E_{1}}&\dots&\delta_{E_{n}}\end{smallmatrix}\right)^{T}.$ Then we arrive at the quadratic minimization problem
(13) $\text{min}_{x^{\prime}_{D}\in\mathbb{R}^{n}}\{{x^{\prime}_{D}}^{T}Kx^{\prime}_{D}-b^{T}x^{\prime}_{D}\},$
where $K$ and $b$ are the coefficient matrix (symmetric) and vector, respectively, extracted from (12). We conclude that the problem of finding the Dirichlet boundary values of the holes reduces to finding a solution to a $n\times n$ linear system of equations. Of course, this construction requires that the necessary bookkeeping of degrees of freedom is available.
4.3 Details of the Reduction Step
The Dirichlet vector $x_{D}$ has a natural partition where the zero Dirichlet part $x_{D_{0}}$ is omitted, with lengths $k_{0},k_{1},\ldots,k_{n}$ such that $m~=~\sum_{j_{0}}^{n}k_{j}$:
$x_{D}=\begin{pmatrix}x_{D_{1}}&x_{E_{1}}&\dots&x_{E_{n}}\end{pmatrix}^{T}.$
The idea is to construct a reduction matrix $R$ which maps, in other words, sums, all coefficients of the dofs $x_{E_{j}}$ to those of $\delta_{E_{j}}$. The matrix $R\in\mathbb{R}^{m\times(n+1)}$ has the form
$R=\begin{pmatrix}\mathbf{1}_{k_{0}}&0_{k_{0}}&\cdots&0_{k_{0}}\\ 0_{k_{1}}&\mathbf{1}_{k_{1}}&&0_{k_{1}}\\ \vdots&&\ddots&\vdots\\ 0_{k_{n}}&0_{k_{n}}&\cdots&\mathbf{1}_{k_{n}}\end{pmatrix}$.
The matrix $A_{DD}$ is block diagonal, thus $K_{0}=R^{T}A_{DD}R$ is a diagonal matrix. The second part $K_{1}=R^{T}A_{DI}A_{II}^{-1}A_{ID}R$ is a symmetric matrix. Then, using the colon notation, we get
$b=K_{1}(1,2:)^{T},\qquad K=K_{1}(2:,2:)-K_{0}(2:,2:).$
The constant term does not affect the minimization.
4.4 Effect of the Improved Construction: Two Holes on a Sphere Revisited
In our earlier work [10] we considered a problem with two holes on a sphere. There the conjugate problem was constructed using our original method that was constrained by the optimization process. Here we first apply the quadratic minimization and, indeed, obtain more accurate results using exactly the same finite element discretization. This is illustrated in Figure 2. What is even more remarkable is that this direct construction has a well-defined a priori computational complexity that is often significantly smaller than that of standard optimization.
The precise description of the problem on the parameter plane is as follows: The computational domain is a unit circle centred at $(1/2,1/2)$, with two circular holes with radius =1/4=1/4 and centres at $(1/4,1/4)$ and $(3/4,3/4)$ denoted as $B_{1}$ and $B_{2}$, respectively. The four corners are selected symmetrically at $z_{i}=(1/2+\cos(\pi(k-1)/2),1/2+\sin(\pi(k-1)/2))$, $k=1,\dots,4$. The optimization step gives us potentials $v_{1}=0.5344459257688663$ on $\partial B_{1}$ and $v_{2}=0.4883797948571419$ on $\partial B_{2}$. The moduli obtained with the new construction of the conjugate problem are
(14) $M(Q)=0.7901907691109106,\;M(\widetilde{Q})=1.2655172762842442,$
with $\operatorname{reci}(Q)=6.987019163240404\times 10^{-8}$.
Figure 2: Two Holes Revisited. Convergence of the reciprocal error as a function of degrees of freedom using optimization and the new construction. The current version is eventually more accurate on exactly the same mesh, and dips below the threshold limit of $10^{-7}$ that indicates the best possible accuracy in the previous version (loglog-plot).
5 Torii
Even though the conjugate function method is not directly applicable on surfaces of genus 1, if the parameter domain can be decomposed into admissible parts, then the conformal map can be constructed by considering each part independently.
5.1 Torus: Genus 1
For a regular torus the parameter domain is simply $0,2\pi]\times[0,2\pi]$. Due to symmetry, one admissible partition is to divide the parameter domain into four parts. The resulting map and the convergence graph are shown in Figure 4. The parameterisation of the torus, with inner radius $r=1/2$, and outer $R=3/2$ is
(15) $T(u,v)=\left(\left(\frac{\cos(u)}{2}+1\right)\cos(v),\left(\frac{\cos(u)}{2}+1\right)\sin(v),\frac{\sin(u)}{2}\right).$
It is known that this parameterisation is not isothermal. This is immediately visible in Figure 4a. For $T(u,v)$, one has
$J_{\mathbf{x}}=\left(\begin{array}[]{cc}-\frac{1}{2}\sin(u)\cos(v)&-\left(\left(\frac{\cos(u)}{2}+1\right)\sin(v)\right)\\ -\frac{1}{2}\sin(u)\sin(v)&\left(\frac{\cos(u)}{2}+1\right)\cos(v)\\ \frac{\cos(u)}{2}&0\\ \end{array}\right),$
$G_{K}^{-1}=\left(\begin{array}[]{cc}4&0\\ 0&\frac{4}{\cos^{2}(u)+4\cos(u)+4}\\ \end{array}\right),$
$\sqrt{\operatorname{det}(G_{K})}=\frac{1}{4}\left(\cos(u)+2\right).$
resulting in
(16) $G_{K}^{-T}G_{S}G_{K}^{-1}\sqrt{\operatorname{det}(G_{K})}=\left(\begin{array}[]{cc}4&0\\ 0&\frac{4}{(\cos(u)+2)^{2}}\\ \end{array}\right)\neq I.$
Indeed, with a suitable reparameterisation of $v$ one can find an isothermal parameterisation for a torus.
Figure 3: Torus with checkerboard colouring of the conformal map.
(a) $(u,v)\in[0,\pi]\times[0,\pi]$
(b) Error estimates vs NN.
Figure 4: Torus. (a) Map on the parameter space. (b) Convergence in the non isothermal parameterisation, error estimates vs. NN (the number of degrees of freedom). Reciprocal error estimate: Solid line.
5.2 Two examples: Genus 2 and Genus 3
Finding conformal mappings on arbitrary surfaces is equivalent to quad meshing on surfaces of arbitrary genus. Euler characteristic cannot be satisfied except for a case of simple torus. In other words, extraordinary points are inevitable in the general case and at best one can find a quasiconformal map. The conjugate function method can be applied to surfaces of higher genus provided that suitable charts can be found. One pair of such realizations is shown in Figure 5. The 2-torus is defined by
(17) $\frac{3}{32}\left(x^{2}+2y^{2}\right)^{2}+\frac{1}{2\left(\left(x-\frac{3}{4}\right)^{2}+y^{2}\right)+\frac{1}{2}}+\frac{1}{2\left(\left(x+\frac{3}{4}\right)^{2}+y^{2}\right)+\frac{1}{2}}+2z^{2}=\frac{13}{10}$
and the 3-torus by
(18) $\frac{1}{24}\left(x^{2}+y^{2}\right)^{2}+\frac{2}{4x^{2}+4(y-2)y+5}+$ $\frac{2}{4x\left(x-\sqrt{3}\right)+4y(y+1)+5}+$
$\hskip-21.52771pt\frac{2}{4x\left(x+\sqrt{3}\right)+4y(y+1)+5}+z^{2}=\frac{13}{10}.$
Figure 5: Double and Triple Tori. Possible realizations.
Let us consider the 2-torus in more detail. What is it that we see in Figure 5? The image is composed of eight sections using symmetries. Let $k_{f}$ be the number of edges of a face $f$ and $\operatorname{val}(v)$ the valence of a vertex $v$. Since
(19) $\sum_{v}\operatorname{val}(v)\;=\;2E\;=\;\sum_{f}k_{f},$
and Euler’s relation $\chi=V-E+F$, one obtains
(20) $\sum_{v}\bigl{(}4-\operatorname{val}(v)\bigr{)}\;-\;\sum_{f}\bigl{(}k_{f}-4\bigr{)}\;=\;4\chi.$
Equivalently,
$\sum_{v}\bigl{(}4-\operatorname{val}(v)\bigr{)}\;=\;4\chi\;+\;\sum_{f}\bigl{(}k_{f}-4\bigr{)}.$
Using standard terminology, the left-hand side is the sum of vertex defects that the extraordinary vertices contribute, and $\sum_{f}(k_{f}-4)$ is the sum of face defects, and specifically a pentagon gives +1+1. For a closed orientable surface of genus $g$,
$\chi\;=\;2-2g.$
With $g=2$, we have $\chi=-2$, so $4\chi=-8$. If all faces are quads ($k_{f}=4$ for all $f$), then $\sum_{f}(k_{f}-4)=0$. Now, (20) gives
$\sum_{v}(4-\operatorname{val}(v))\;=\;-8,$
so we need for a valid subdivision of the surface, for example, eight pentagons. By close inspection of the checkerboard image of the 2-torus, one can see that there is, indeed, one pentagon per every symmetric section, and hence, we conclude that the quasiconformal map is valid.
6 Numerical Experiments on Domains with Multiple Boundary Components
The efficacy of the conjugate function method on surfaces has been established already before [10]. In this section the focus is on problems with multiple boundary components both on planar domains and surfaces. The first experiment was originally proposed by Nasser [20]. The latter two are variations of the same theme: slit domains with strong singularities. First a selection of slits in a rectangle is considered and then lifted onto a hemisphere that passes through the corner points. All geometric specifications are available from the authors upon request. The three cases are outlined in Table 1.
Table 1: Description of the three numerical experiments. The number of elements includes both triangles and quadrilaterals. The polynomial order $p$ is the maximal used in the experiments and the number of degrees of freedom corresponds to this value. In Cases B and C, the mesh used is exactly the same.
Case Surface #holes #(nodes,edges,elements) $p$ # dof
A plane 35 (23336,64018,40639) 8 2,062,192
B plane 50 (8931,17080,8100) 10 1,163,711
C hemisphere 50 (8931,17080,8100) 10 1,163,711
6.1 Case A: Nasser’s Challenge
In this example 35 holes are punched inside a disk. The holes are carefully selected to include either, one, two, or four reentrant corners or cusps.
Together with the map on the domain (Figure 6) and the canonical domain (Figure 7a) one can trace the paths from edge to edge where the gridlines do or do not touch any of the holes.
Figure 6: Nasser’s Configuration: Map
The convergence graphs are shown in Figure 7c. Since the mesh (see Figure 7b) is generated on the original boundary discretisation, we cannot expect exponential convergence. Indeed, even though the reciprocal error is relatively small, the observed rate is only algebraic. The auxiliary subspace estimates for the primary and conjugate problems are optimistic. Notice that since the reciprocal error measures a product, it should stay above the auxiliary subspace estimates.
(a) Canonical domain.
(b) Finite element mesh.
(c) Convergence: Error estimates vs NN.
Figure 7: Nasser’s Configuration. (a) Canonical domain. (b) Mesh used in computations. (c) Convergence graphs for error estimates. NN is the number of degrees of freedom. The data points represent the results on a fixed mesh with the constant polynomial order ranging from $p=2,\ldots,10$. Solid line: Reciprocal error; Dashed and dotted: Auxiliary subspace estimates for the primary and conjugate problems, respectively (loglog-plot).
6.2 Case B: Random Segments Inside a Rectangle
As discussed above, with proper grading of the meshes, the $hp$-FEM should converge exponentially even when strong singularities are present. Here 50 slits are placed at random in a $40\times 40$ rectangle resulting in 100 $2\pi$-corners inside the domain.
The resulting map is shown in Figure 8. The detail shown in Figure 8b shown how the slit perturbs the ”flow lines” locally. The slit canonical domain is illustrated in Figure 9a. Together with the map on the domain (see Figure 8a) one can trace the paths from edge to edge where the gridlines do or do not touch any of the slits.
(a) Conformal map on the domain.
(b) Detail of the map.
Figure 8: Random Segments. Conformal map with a detail.
The convergence graphs are shown in Figure 9b. On a graded mesh with a constant polynomial order, we observe exponential convergence in the reciprocal error. The auxiliary subspace estimates for the primary and conjugate problems exhibit the same rate giving us high confidence in the results.
(a) Canonical domain.
(b) Convergence: Error estimates vs NN.
Figure 9: Random Segments. (a) Canonical slit domain. (b) Convergence graphs for error estimates. NN is the number of degrees of freedom. The data points represent the results on a fixed mesh with the constant polynomial order ranging from $p=2,\ldots,10$. Solid line: Reciprocal error; Dashed and dotted: Auxiliary subspace estimates for the primary and conjugate problems, respectively (loglog-plot).
6.3 Case C: Random Segments on a Hemisphere
We continue with the same planar configuration. However, now the configuration represents the parameter space that is lifted onto a conforming hemisphere (Figure 10). Interestingly, the convergence behaviour of the $hp$-FEM is practically unchanged despite the added complexity of the Laplace-Beltrami problem (Figure 11c). This result underlines the fact that the method for constructing the conjugate problem depends only on the discretisation and only indirectly on the underlying variational formulation.
Of particular interest is the comparison of the two canonical domains shown in Figure 11b. If one considers the two configurations, planar and surface, as extremal configurations, it is clear that the intermediate stages in continuous lifting would result in smooth transition from one to another. This immediately suggests applications based on transformations on the canonical domain.
(a) Conformal map on the parameter domain.
(b) Conformal map on the domain.
Figure 10: Random Segments on Hemisphere.
(a) Canonical domain.
(b) Comparison of the canonical domains.
(c) Convergence: Error estimates vs NN.
Figure 11: Random Segments on Hemisphere. (a) Canonical slit domain. (b) Comparison of the canonical domains: Black for planar and red for the surface. (c) Convergence graphs for error estimates. NN is the number of degrees of freedom. The data points represent the results on a fixed mesh with the constant polynomial order ranging from $p=2,\ldots,10$. Solid line: Reciprocal error; Dashed and dotted: Auxiliary space estimates for the primary and conjugate problems, respectively (loglog-plot).
6.4 Comment on Computational Complexity
The timing data over the random segment case is shown in Figure 12. Even in the implementation without a native sparse Cholesky decomposition, the construction of the conjugate problem is roughly comparable to one assembly or integration of the stiffness matrix. In (12) the latter term could be evaluated more efficiently if the Cholesky decomposition $L_{II}$ of the matrix $A_{II}=L_{II}L_{II}^{T}$ were available.
Figure 12: Timing information on Random Segments. The data points represent the results on a fixed mesh with the constant polynomial order ranging from $p=2,\ldots,10$. Solid line: Integration; Dashed: Construction of the conjugate problem (loglog-plot). The lack of sparse Cholesky decomposition leads to a bottleneck in (12).
7 Application: Texture Mapping
In this section we bring together the many aspects considered above. We use a simplified image of a Chinese opera mask as the model [^1] This image is derived from the original by Yudhi Sholihana, https://www.vecteezy.com/free-vector/peking-opera. In Figure 13 this image is lifted onto a surface and the conformal map is laid onto it.
Since in this example there are two holes (only for the eyes), the canonical domain in Figure 14a is relatively simple. However, the high accuracy of the method is reflected in the mapped image in Figure 14b. The intricate details of the design are clearly visible. Combined with the observation on the two canonical domains above, this strongly suggests that these methods could be used in animation as suggested in [6].
(a) Original image.
(b) Conformal map on the parameter domain.
(c) Conformal map on the domain.
Figure 13: Opera Mask. Conformal map on the domain.
(a) Canonical domain.
(b) Image on the canonical domain.
Figure 14: Opera Mask. Image of the mask on the canonical domain. The slits are indicated with green lines.
8 Conclusions
In this work we have generalised and refined the conjugate function method to achieve the same level of accuracy on simply and multiply connected planar domains and Riemann surfaces. In contrast to the Koebe iteration, the construction presented here can be viewed as direct, with predictable a priori computational complexity. Our implementation relies on high-order finite element methods, both for obtaining high accuracy and the construction of the conjugate problem. This does not mean that the key observation in the minimisation process depends on one particular method for solving the underlying partial differential equations.
Extension to high-genus surfaces remains a future challenge. Automatic detection of charts for an atlas is a very difficult problem and we have not made any attempts to address this question. In cases where the symmetries suggest parameterisations, such as the 2-torus and 3-torus considered above, the conjugate function method can be used to compute highly accurate quasiconformal maps.
Acknowledgments
We thank warmly prof. M.M.S. Nasser for kindly providing us with one of the geometric configurations used in one of the experiments.
H. Hakula was supported by the Research Council of Finland (Flagship of Advanced Mathematics for Sensing Imaging and Modelling grant 359181).
A. Rasila and Y. Zheng were supported by NSF of Guangdong Province (No. 2024A1515010467) and Li Ka Shing Foundation GTIIT-STU Joint Research Grant (No. 2024LKSFG06)

### v2

@@
 The conjugate function method is an algorithm for numerical computation of conformal mappings for simply and multiply connected domains on surfaces.
-In this paper the conjugate function method is generalized and refined to achieve the same level of accuracy on simply and multiply connected planar domains and Riemann surfaces.
-The main challenge addressed here is the accurate and efficient construction of the conjugate problem for multiply connected domains.
+In this paper the conjugate function method, earlier used for simply connected domains, is generalized and refined to achieve the same level of accuracy on multiply connected planar domains and Riemann surfaces.
+The main challenge is the accurate and efficient construction of boundary values for the conjugate problem on multiply connected domains.
 The method relies on high-order finite element methods which allow for highly accurate computations of mappings on surfaces, including domains of complex boundary geometry containing strong singularities and cusps.
+We also derive the reciprocal error estimate for the multiply connected case.
 The efficacy of the proposed method is illustrated via an extensive set of numerical experiments with error estimates.
@@
 {AMS}
-30Cxx, 30Fxx, 65E10, 65N30
+30C85, 30F10, 31A15, 65E05, 65E10, 65N30
 1 Introduction
@@
 [5]) and mathematical geodesy (see [27]).
-Traditionally, these applications have been mostly restricted on conformal mappings between planar domains.
-Even in this setting finding mappings in the case of multiply connected domains has remained challenging.
+Traditionally, numerical methods of approximating conformal mappings have been mostly restricted to planar domains.
+Even in this setting finding mappings between multiply connected domains has remained challenging.
 M.M.S.
 Nasser has used the boundary integral equation (BIE) with impressive results [19, 21].
-Moreover, in a very recent work, Nasser discusses how to implement the generalized Koebe’s iterative method efficiently [15].
+Furthermore, in a very recent work, Nasser discusses efficient implementation of the generalized Koebe’s iterative method [15].
 Koebe’s iterative method can also be implemented for surfaces.
@@
 Due to both theoretical and implementation complexity, there are relatively few publications addressing specifics.
-See, for instance, [16, 17, 18, 25, 28].
+See, for example, [16, 17, 18, 25, 29].
 Even though a properly implemented Koebe’s iteration converges reasonably fast (at least in the planar case), it is very difficult to provide any reliable error estimates or convergence rates.
 Our method of choice for conformal mappings, the conjugate function method, was first introduced in [8] for simply and doubly connected planar domains.
-We use the $hp$-adaptive finite element method ($hp$-FEM) introduced in [11] for computing potential functions and conformal moduli, but other methods can be used as well, e.g., [13] the stochastic version of the algorithm was given for circular arc domains in the plane.
-One advantage of $hp$-FEM is that in error analysis than not only follows from the adoption of the finite element method for implementation but is also inherent in the form of the reciprocal identity relating the conformal capacities of the primary and conjugate problems.
-In this work this concept is made precise for surfaces in the specific setting of Q-type canonical domains.
+We use the $hp$-adaptive finite element method ($hp$-FEM) introduced in [11] to compute potential functions and conformal moduli, but other methods can also be used, e.g., in [13] the stochastic version of the conjugate function algorithm was given for circular arc domains in the plane.
+An advantage of $hp$-FEM is the mature error analysis.
+Interestingly, the reciprocal identity relating the conformal capacities of the primary and conjugate problems can also be viewed as an error indicator.
+In this work this concept is made precise for surfaces with multiple boundary components in the specific setting of so-called Q-type canonical domains (see [9] and below).
 On method development the main contribution of this work is the new, direct algorithm for the construction of the conjugate problem for multiply connected problems.
@@
 Not only is the approach much faster than the previous ones, but the accuracy of the solution of the conjugate problem is brought to the same level as that of the primary problem.
-The same algorithm applies to problems on surfaces as well without any modifications.
-The conjugate function method is now comparable to the state-of-the-art methods with the additional benefits of the high-order finite element frameworks such as advanced error estimation and adaptivity.
+The same algorithm applies to problems on multiply connected domains on surfaces as well without any modifications.
 1.1 Illustrative Example: 3D Face Recognition
@@
 Since a standard 2D image recognition is fragile due to many factors, 3D methods have been proposed as a more reliable alternative.
-Each 3D surface with suitable topology can be mapped conformally to a 2D domain and thus, the image recognition becomes a conformal map comparison problem.
+Each 3D surface with suitable topology can be mapped conformally to a 2D domain, and thus, the image recognition becomes a conformal map comparison problem.
 See [26].
@@
 First, four points on the outer boundary are fixed.
-In our formulation the canonical domain is a quadrilateral with horizontal slits if holes are present.
+In our formulation, the canonical domain is a quadrilateral with horizontal slits if holes are present.
 Second, the Laplace-Beltrami equation is solved twice with different boundary conditions.
 As mentioned above, this process is only marginally slower if multiple holes are present.
-In the example both eyes and the mouth are modelled as holes.
+In the example, both eyes and the mouth are modeled as holes.
 Figure 1: Map of a face. Left: The map on the parameter space. Right: Checkerboard illustration of the map on the surface. Original data source: [24].
 1.2 Organization
-The rest of the paper is organized as follows: After the introduction, preliminaries including the finite element method used and the planar conjugate function method are covered.
+The rest of the paper is organized as follows: After the introduction, preliminaries including the finite element method used are presented.
+In Section 3 the conjugate function method is introduced.
+Section 4 discusses both the proof of the reciprocal identity for oriented surfaces, and the efficient construction of the conjugate problem.
+Torii are given as a special class of closed surfaces in Section 5.
+Sections 6-8 cover series of numerical experiments and applications, before conclusions at the end.
 2 Preliminaries
-In section some basic concepts and definitions used throughout this paper are introduced.
+In this section, some basic concepts and definitions used throughout this paper are introduced.
 2.1 Riemann surfaces
-Recall that a Hausdorff space $X$ is locally Euclidean of dimension $n\geq 1$ if every point $p$ in $X$ has a neighbourhood $U$ such that there is a homeomorphism $\varphi$ from $U$ onto an open subset of $\mathbb{R}^{n}$.
-We call the pair $(U,\ \varphi:U\to\mathbb{R}^{n})$ a chart, $U$ a coordinate neighbourhood or a coordinate open set, and $\phi$ a coordinate mapping or a coordinate system on $U$.
-We say that a chart $(U,\varphi)$ is a centered at $p\in U$ if $\varphi(p)=0$.
-Let $X$ be a two-dimensional manifold, which we also consider as a one-dimensional complex manifold.
-A complex chart on $X$ is a homeomorphisms $\varphi:U\to V$ from an open subset $U\subset X$ to an open subset $V\subset\mathbb{C}$.
-Two complex charts $\varphi_{j}:U_{j}\to V_{j}$ (for $j=1,2$) are said to be holomorphically compatible if the mapping
+Let $X$ be a Hausdorff space that is locally homeomorphic to the Euclidean plane $\mathbb{R}^{2}$, which we also consider as a one-dimensional complex manifold.
+We assume that its boundary $\Gamma$ is either empty or consists of a finite number of rectifiable arcs that may intersect only at their endpoints, and its closure $X\cup\Gamma$ is compact.
+A complex chart on $X$ is a homeomorphism $\varphi:U\to V$ from an open subset $U\subset X$ onto an open subset $V\subset\mathbb{C}$.
+Two complex charts $\varphi_{j}:U_{j}\to V_{j}$ (for $j=1,2$) are said to be holomorphically compatible if $U_{1}\cap U_{2}=\emptyset$ or the mapping
 $\varphi_{2}\circ\varphi_{1}^{-1}:\varphi_{1}(U_{1}\cap U_{2})\to\varphi_{2}(U_{1}\cap U_{2})$
@@
 Intuitively, this corresponds to the number of holes on the surface.
-Note that in this paper we consider both closed surfaces and surfaces with multiple boundary components.
-For more information of Riemann surfaces, we refer to [4].
+Note that in this paper, we consider both closed surfaces and surfaces with multiple boundary components.
+For more information about Riemann surfaces, we refer to [4, 14].
 2.2 High-Order Finite Element Method
-High-order finite element methods have the capability for exponential convergence provided the discretization is constructed properly in both domain and in local polynomial order ($hp$-version).
-In many cases one has to work with cases where exact parameterization of the boundaries on the parameter space is not known.
+High-order finite element methods have the capability for exponential convergence in conformal capacity problems provided that for a given discretization the polynomial order (hp-version) is properly chosen and the discretization is correctly adjusted to reflect the intricacies of the domain boundary.
+In many cases one has to work with cases where the exact parameterization of the boundaries on the parameter space is not known.
 Then only algebraic convergence rate is to be expected.
@@
 2.2.1 Error Estimation
-The solution method specific a posteriori error estimation method applied here is the so-called auxiliary subspace error estimation.
+The a posteriori error estimation method specific to the solution method applied here is the so-called auxiliary subspace error estimation.
 The estimate measures the projection of the residual to the auxiliary space, that is, to a set of finite element degrees of freedom that extend the approximation space in a natural way.
@@
 Our approach for geometric grading of the meshes with known locations of singularities is based on rule based algorithms [12].
-One non-standard boundary type in the context of finite element method considered here is the slit or cut.
-We define a cut to be a set of element nodes and edges that form a loop with zero area.
+One non-standard boundary type in the context of finite element method considered here is the slit.
+We define a slit to be a set of element nodes and edges that form a loop with zero area.
 That is, the data structures support having two edges on top of each other, both with their own kinematic constraints.
 3 The Conjugate Function Method On Planar Domains and Riemann Surfaces
-Let $\Omega\subset\mathcal{S}$ is a simply connected domain on a Riemann surface $\mathcal{S}$ so that its boundary of $\Omega$ is a Jordan curve.
+Let $\Omega\subset\mathcal{S}$ be a simply connected domain on a Riemann surface $\mathcal{S}$ so that the boundary of $\Omega$ is a Jordan curve.
 We call $\Omega$ together with four positively oriented points $z_{1},z_{2},z_{3},z_{4}\in\partial\Omega$ a (generalized) quadrilateral and denote it by $Q=(\Omega;z_{1},z_{2},z_{3},z_{4})$.
@@
 where $n$ is the unit exterior boundary normal and $\Delta_{S}$ is the Laplace-Beltrami operator.
-Then the conformal modulus is connected to the above boundary value problem by the identity (see e.g.
+For $\Omega\subset\mathbb{C}$ the conformal modulus is connected to the above boundary value problem by the identity (see e.g.
 Ahlfors [1, Theorem 4.5] and Papamichael and Stylianopoulos [22, Theorem 2.3.3]):
 (2) $M(Q)=\iint_{\Omega}|\nabla u|^{2}\,dx\,dy.$
-For a quadrilateral $Q=(\Omega;z_{1},z_{2},z_{3},z_{4})$ we call $\widetilde{Q}=(\Omega;z_{2},z_{3},z_{4},z_{1})$ its conjugate quadrilateral and the corresponding problem (1) for quadrilateral $\widetilde{Q}$ the conjugate Dirichlet–Neumann problem.
-It is well-known that if $M(Q)=h>0$, then $M(\widetilde{Q})=1/h$, which leads to the following very useful reciprocal identity: For all quadrilaterals $Q$, we have
-(3) $M(Q)M(\widetilde{Q})=1.$
-Furthermore, we may observe that canonical conformal mapping of a quadrilateral $Q=(\Omega;z_{1},z_{2},z_{3},z_{4})$ onto the rectangle $R_{h}$ with vertices at $1+ih$, ihih, 0, and $1$, can be obtained by solving the corresponding Dirichlet–Neumann problem $u$ and its conjugate problem.
+For a quadrilateral $Q=(\Omega;z_{1},z_{2},z_{3},z_{4})$ we call $\widetilde{Q}=(\Omega;z_{2},z_{3},z_{4},z_{1})$ its conjugate quadrilateral and the corresponding problem (1) for the quadrilateral $\widetilde{Q}$ the conjugate Dirichlet–Neumann problem.
+It is well-known that if $M(Q)=h>0$, then $M(\widetilde{Q})=1/h$, which leads to
+(3) $M(Q)M(\widetilde{Q})=1,$
+for all quadrilaterals $Q$. This reciprocal identity is very useful since it can be interpreted as an error estimate [11].
+Furthermore, we may observe that the canonical conformal mapping of a quadrilateral $Q=(\Omega;z_{1},z_{2},z_{3},z_{4})$ onto the rectangle $R_{h}$ with vertices at $1+ih$, ihih, 0, and $1$, can be obtained by solving the corresponding Dirichlet–Neumann problem and its conjugate problem.
 Lemma 3.1.
 Let $Q$ be a quadrilateral with modulus $h$, and suppose $u$ solves the Dirichlet–Neumann problem (1).
-If $v$ is a harmonic function conjugate to $u$, satisfying $v(\operatorname{Re}z_{3},\operatorname{Im}z_{3})=0$, and $\tilde{u}$ represents the harmonic solution for the conjugate quadrilateral $\widetilde{Q}$, then $v=h\tilde{u}$.
+If $v$ is a harmonic function conjugate to $u$, satisfying $v(\operatorname{Re}z_{3},\operatorname{Im}z_{3})=0$, and $\tilde{u}$ represents the harmonic function solving the Dirichlet–Neumann problem for the conjugate quadrilateral $\widetilde{Q}$, then $v=h\tilde{u}$.
 3.1 Q-type canonical domains
-Let $h>0$, and let $\big{(}(\zeta_{1},d_{1}),\ldots,(\eta_{N},d_{N})\big{)}$ be complex numbers such that $\operatorname{Re}\zeta_{i}\in(0,1)$ and $\operatorname{Im}\zeta_{i}\in(0,h)$ for all $i=1,\ldots,N$, where $N\geq 1$.
-For a multiply connected plane case, we will consider canonical domains of the type
-(4) $\big{(}(0,1)\times(0,h)\big{)}\setminus\bigcup_{j=1}^{N}[\zeta_{j},\zeta_{j}+d_{j}],$
+Let $h>0$, and let $\big((\zeta_{1},d_{1}),\ldots,(\zeta_{N},d_{N})\big)$ be complex numbers such that $\operatorname{Re}\zeta_{i}\in(0,1)$ and $\operatorname{Im}\zeta_{i}\in(0,h)$ for all $i=1,\ldots,N$, where $N\geq 1$.
+For the case of multiply connected planar domains, we will consider canonical domains of the type
+(4) $\big((0,1)\times(0,h)\big)\setminus\bigcup_{j=1}^{N}[\zeta_{j},\zeta_{j}+d_{j}]\subset\mathbb{C},$
 i.e., a rectangle $(1,0)\times(0,h)$ with $N$ horizontal slits removed. Canonical domains of this type are called quadrilateral-like (or Q-type) in [9].
@@
 Proposition 4.1.
-Let $\Omega$ be a domain on an oriented surface $S$ that can mapped conformally onto a Q-type canonical domain $Q_{*}=(\Omega;z_{1},z_{2},z_{3},z_{4})$ defined by (4), where the points $z_{1},z_{2},z_{3},z_{4}$ are chosen from the same boundary component of $\Omega$.
+Let $\Omega$ be a domain on an oriented surface $S$ that can be mapped conformally onto a Q-type canonical domain $Q_{*}=(\Omega;z_{1},z_{2},z_{3},z_{4})$ defined by (4), where the points $z_{1},z_{2},z_{3},z_{4}$ are chosen from the same boundary component of $\Omega$.
 Then the generalized conformal moduli of $\Omega$ satisfy the reciprocal identity:
@@
 But in this case, the solution of the boundary value problem (8) is given by the harmonic functions $u(x,y)=x$ and $v(x,y)=y/h$. Therefore, the desired identity follows immediately.
+Remark 4.3.
+Previously, the use of reciprocal identity as a method of error estimation has been verified only in the simply connected case [8].
 4.2 Linear Algebra Based Construction of the Conjugate Problem
-Let us divide the degrees of freedom, internal and boundary, into sets.
-The sets of a discretized quadrilateral are denoted $B,D_{0},D_{1},N^{0}$, and $N^{1}$, for internal, Dirichlet $u=0$, Dirichlet $u=1$, Neumann with Dirichlet $u=0$ in the primary problem, and Neumann with Dirichlet $u=1$ in the primary problem, degrees of freedom, respectively.
-The degrees of freedom on the boundaries of the holes $E_{i}$ are all Dirichlet, constant over the boundary and initially unknown $\delta_{E_{i}}$.
-The discretized system is
+As indicated in (8), the discretized primary and conjugate problems share the same degrees of freedom before application of boundary conditions.
+The degrees of freedom admit a partition that can be used to define a block structure for the discretized matrix $A$.
+For the conjugate problem the degrees of freedom on the four boundary segments $\gamma_{i}$ defining the quadrilateral $Q$ are denoted by $D_{0},D_{1},N^{0}$, and $N^{1}$, where $D_{0}$ indicates $v=0$ and $N^{0}$ zero Neumann on the segment with $u=0$ on the primary problem.
+The internal degrees of freedom are denoted by $B$.
+Finally, the degrees of freedom on the boundaries of the holes are denoted by $E_{i}$ .
+Remark 4.4.
+Since the potential is constant over the boundary of every hole in the conjugate problem, all edge degrees of freedom are set to zero.
+This is a technical detail that does not affect the discussion below.
+Using the notation given above the discretized system has the following block structure
 $A=\begin{pmatrix}A_{BB}&A_{BN^{1}}&A_{BN^{0}}&A_{BD_{1}}&A_{BD_{0}}&A_{BE_{1}}&\dots&A_{BE_{n}}\\ A_{N^{1}B}&A_{N^{1}N^{1}}&A_{N^{1}N^{0}}&A_{N^{1}D_{1}}&A_{N^{1}D_{0}}&A_{N^{1}E_{1}}&&A_{N^{1}E_{n}}\\ A_{N^{0}B}&A_{N^{0}N^{1}}&A_{N^{0}N^{0}}&A_{N^{0}D_{1}}&A_{N^{0}D_{0}}&A_{N^{0}E_{1}}&&A_{N^{0}E_{n}}\\ A_{D_{1}B}&A_{D_{1}N^{1}}&A_{D_{1}N^{0}}&A_{D_{1}D_{1}}&A_{D_{1}D_{0}}&A_{D_{1}E_{1}}&&A_{D_{1}E_{n}}\\ A_{D_{0}B}&A_{D_{0}N^{1}}&A_{D_{0}N^{0}}&A_{D_{0}D_{1}}&A_{D_{0}D_{0}}&A_{D_{0}E_{1}}&&A_{D_{0}E_{n}}\\ A_{E_{1}B}&A_{E_{1}N^{1}}&A_{E_{1}N^{0}}&A_{E_{1}D_{1}}&A_{E_{1}D_{0}}&A_{E_{1}E_{1}}&&A_{E_{1}E_{n}}\\ \vdots&&&&&&\ddots\\ A_{E_{n}B}&A_{E_{n}N^{1}}&A_{E_{n}N^{0}}&A_{E_{n}D_{1}}&A_{E_{n}D_{0}}&A_{E_{n}E_{1}}&\dots&A_{E_{n}E_{n}}\\ \end{pmatrix}$.
@@
 $x=\begin{pmatrix}x_{B}&x_{N^{1}}&x_{N^{0}}&x_{D_{1}}&x_{D_{0}}&x_{E_{1}}&\dots&x_{E_{n}}\end{pmatrix}^{T}.$
-In order to derive the quadratic form we first divide the degrees of freedom into free, that is, zero Neumann, and fixed, where the internal and Neumann degrees of freedom are free, and all Dirichlet ones are fixed.
-These sets are denoted by $I$ and $D$, respectively.
-The matrix $A$ has a simple $2\times 2$ form $A=\left(\begin{smallmatrix}A_{II}&A_{ID}\\ A_{DI}&A_{DD}\\ \end{smallmatrix}\right)$, and $x=\left(\begin{smallmatrix}x_{I}&x_{D}\end{smallmatrix}\right)^{T}$.
+In order to derive the quadratic form we further divide the degrees of freedom into two classes: free ($B$, $N^{0}$, $N^{1}$) and fixed (all Dirichlet).
+These classes are denoted by $I$ and $D$, respectively.
+Then the matrix $A$ has a simple $2\times 2$ form $A=\left(\begin{smallmatrix}A_{II}&A_{ID}\\ A_{DI}&A_{DD}\\ \end{smallmatrix}\right)$, and $x=\left(\begin{smallmatrix}x_{I}&x_{D}\end{smallmatrix}\right)^{T}$.
 Using the block form one can solve (formally) the free degrees of freedom
 $A_{II}x_{I}=-A_{ID}x_{D},\quad x_{I}=-A_{II}^{-1}A_{ID}x_{D}.$
-Using this, one can then evaluate the quadratic form
+Therefore, one can simplify the quadratic form
 (12) $Q_{F}=\begin{pmatrix}x_{I}\\ x_{D}\\ \end{pmatrix}^{T}\begin{pmatrix}A_{II}&A_{ID}\\ A_{DI}&A_{DD}\\ \end{pmatrix}\begin{pmatrix}x_{I}\\ x_{D}\\ \end{pmatrix}=x_{D}^{T}A_{DD}x_{D}-x_{D}^{T}A_{DI}A_{II}^{-1}A_{ID}x_{D}.$
-Let us next denote the vector of unknown parameter values $x^{\prime}_{D}=\left(\begin{smallmatrix}\delta_{E_{1}}&\dots&\delta_{E_{n}}\end{smallmatrix}\right)^{T}.$ Then we arrive at the quadratic minimization problem
+Let us next denote the vector of unknown parameter values $x^{\prime}_{D}=\left(\begin{smallmatrix}v_{1}&\dots&v_{n}\end{smallmatrix}\right)^{T}.$ Using the fact that every nonzero degree of freedom on $E_{i}$ is constant and equal to $v_{i}$, we arrive at the quadratic minimization problem
 (13) $\text{min}_{x^{\prime}_{D}\in\mathbb{R}^{n}}\{{x^{\prime}_{D}}^{T}Kx^{\prime}_{D}-b^{T}x^{\prime}_{D}\},$
 where $K$ and $b$ are the coefficient matrix (symmetric) and vector, respectively, extracted from (12).
-We conclude that the problem of finding the Dirichlet boundary values of the holes reduces to finding a solution to a $n\times n$ linear system of equations.
+We conclude that the problem of finding the Dirichlet boundary values of the holes reduces to finding a solution to an $n\times n$ linear system of equations.
 Of course, this construction requires that the necessary bookkeeping of degrees of freedom is available.
@@
 $x_{D}=\begin{pmatrix}x_{D_{1}}&x_{E_{1}}&\dots&x_{E_{n}}\end{pmatrix}^{T}.$
-The idea is to construct a reduction matrix $R$ which maps, in other words, sums, all coefficients of the dofs $x_{E_{j}}$ to those of $\delta_{E_{j}}$.
-The matrix $R\in\mathbb{R}^{m\times(n+1)}$ has the form
+The idea is to construct a reduction matrix $R$ which maps, in other words, sums, all coefficients of the dofs $x_{E_{j}}$ to those of $v_{j}$. The matrix $R\in\mathbb{R}^{m\times(n+1)}$ has the form
 $R=\begin{pmatrix}\mathbf{1}_{k_{0}}&0_{k_{0}}&\cdots&0_{k_{0}}\\ 0_{k_{1}}&\mathbf{1}_{k_{1}}&&0_{k_{1}}\\ \vdots&&\ddots&\vdots\\ 0_{k_{n}}&0_{k_{n}}&\cdots&\mathbf{1}_{k_{n}}\end{pmatrix}$.
@@
 What is even more remarkable is that this direct construction has a well-defined a priori computational complexity that is often significantly smaller than that of standard optimization.
-The precise description of the problem on the parameter plane is as follows: The computational domain is a unit circle centred at $(1/2,1/2)$, with two circular holes with radius =1/4=1/4 and centres at $(1/4,1/4)$ and $(3/4,3/4)$ denoted as $B_{1}$ and $B_{2}$, respectively.
-The four corners are selected symmetrically at $z_{i}=(1/2+\cos(\pi(k-1)/2),1/2+\sin(\pi(k-1)/2))$, $k=1,\dots,4$.
-The optimization step gives us potentials $v_{1}=0.5344459257688663$ on $\partial B_{1}$ and $v_{2}=0.4883797948571419$ on $\partial B_{2}$.
-The moduli obtained with the new construction of the conjugate problem are
-(14) $M(Q)=0.7901907691109106,\;M(\widetilde{Q})=1.2655172762842442,$
-with $\operatorname{reci}(Q)=6.987019163240404\times 10^{-8}$.
+The precise description of the problem on the parameter plane is as indicated in Table 1. The moduli obtained with the new construction of the conjugate problem are
+(14) $M(Q)=0.7901908,\;M(\widetilde{Q})=1.2655173,$
+with $\operatorname{reci}(Q)=7\times 10^{-8}$.
+Table 1: Two Holes: Unit circle centered at $(1/2,1/2)$. The locations, radii, and the obtained potential values in the construction of the conjugate problem are given for both holes.
+Hole Center Radius $v_{k}$ on $\partial B_{k}$
+$B_{1}$ $(1/4,1/4)$ $1/4$ $0.5344459257688663$
+$B_{2}$ $(3/4,3/4)$ $1/4$ $0.4883797948571419$
 Figure 2: Two Holes Revisited.
 Convergence of the reciprocal error as a function of degrees of freedom using optimization and the new construction.
-The current version is eventually more accurate on exactly the same mesh, and dips below the threshold limit of $10^{-7}$ that indicates the best possible accuracy in the previous version (loglog-plot).
+The current version is eventually more accurate on exactly the same mesh, and dips below the threshold limit of $10^{-7}$ that indicates the best possible accuracy when the black-box optimization routine was used (loglog-plot) [10].
 5 Torii
-Even though the conjugate function method is not directly applicable on surfaces of genus 1, if the parameter domain can be decomposed into admissible parts, then the conformal map can be constructed by considering each part independently.
+The conjugate function method is not directly applicable on closed surfaces of genus 1 or higher.
+However, if the parameter domain can be decomposed into admissible parts, then the conformal map can be constructed by considering each part independently.
 5.1 Torus: Genus 1
@@
 The resulting map and the convergence graph are shown in Figure 4.
-The parameterisation of the torus, with inner radius $r=1/2$, and outer $R=3/2$ is
+The parameterisation of the torus, with inner radius $r=1/2$, and outer radius $R=3/2$ is
 (15) $T(u,v)=\left(\left(\frac{\cos(u)}{2}+1\right)\cos(v),\left(\frac{\cos(u)}{2}+1\right)\sin(v),\frac{\sin(u)}{2}\right).$
@@
 (16) $G_{K}^{-T}G_{S}G_{K}^{-1}\sqrt{\operatorname{det}(G_{K})}=\left(\begin{array}[]{cc}4&0\\ 0&\frac{4}{(\cos(u)+2)^{2}}\\ \end{array}\right)\neq I.$
-Indeed, with a suitable reparameterisation of $v$ one can find an isothermal parameterisation for a torus.
+Indeed, with a suitable parameterization of $v$ one can find an isothermal parameterisation for a torus.
 Figure 3: Torus with checkerboard colouring of the conformal map.
@@
 (19) $\sum_{v}\operatorname{val}(v)\;=\;2E\;=\;\sum_{f}k_{f},$
-and Euler’s relation $\chi=V-E+F$, one obtains
+and Euler’s relation states that $\chi=V-E+F$, one obtains
 (20) $\sum_{v}\bigl(4-\operatorname{val}(v)\bigr)\;-\;\sum_{f}\bigl(k_{f}-4\bigr)\;=\;4\chi.$
@@
 so we need for a valid subdivision of the surface, for example, eight pentagons.
-By close inspection of the checkerboard image of the 2-torus, one can see that there is, indeed, one pentagon per every symmetric section, and hence, we conclude that the quasiconformal map is valid.
+By close inspection of the checkerboard image of the 2-torus, one can see that there is, indeed, one pentagon per every symmetric section, and hence, we conclude that the construction is valid.
 6 Numerical Experiments on Domains with Multiple Boundary Components
@@
 The auxiliary subspace estimates for the primary and conjugate problems are optimistic.
-Notice that since the reciprocal error measures a product, it should stay above the auxiliary subspace estimates.
+Notice that since the reciprocal error measures a product, it should stay above the auxiliary subspace estimates of both the primary and the conjugate problems.
 (a) Canonical domain.
@@
 (c) Convergence graphs for error estimates.
-NN is the number of degrees of freedom.
+NN is the number of degrees of freedom that depends on the mesh and the polynomial order.
 The data points represent the results on a fixed mesh with the constant polynomial order ranging from $p=2,\ldots,10$.
@@
 The resulting map is shown in Figure 8.
-The detail shown in Figure 8b shown how the slit perturbs the ”flow lines” locally.
+The detail shown in Figure 8b shown how the slit perturbs the “flow lines” locally.
 The slit canonical domain is illustrated in Figure 9a.
@@
 (b) Convergence graphs for error estimates.
-NN is the number of degrees of freedom.
+NN is the number of degrees of freedom that depends on the mesh and the polynomial order.
 The data points represent the results on a fixed mesh with the constant polynomial order ranging from $p=2,\ldots,10$.
@@
 Figure 14: Opera Mask. Image of the mask on the canonical domain. The slits are indicated with green lines.
+8 Polyhedral Surfaces
+We conclude the set of examples with a polyhedral closed surface [28].
+This is illustrated in Figure 15.
+The original geometric model is triangulated into 5026 3D triangles.
+The primary and conjugate problems are set simply by dividing the bottom rim into four sections.
+The image clearly shows the crowding effect, the sizes of the images of the cells of the uniform grid on the canonical are different on different parts of the domain.
+Since there is no global parameterization of the surface, in other words, no analytic chart is available, the variational formulation is adjusted on every surface triangle.
+Figure 15: Polyhedral surface. The geometric model on the left is triangulated into 5026 3D triangles. Using the bottom of the statue as the boundary, a conformal map of the surface is computed.
 9 Conclusions
-In this work we have generalised and refined the conjugate function method to achieve the same level of accuracy on simply and multiply connected planar domains and Riemann surfaces.
+In this work we have generalized and refined the conjugate function method to achieve the same level of accuracy on simply and multiply connected planar domains and Riemann surfaces.
 In contrast to the Koebe iteration, the construction presented here can be viewed as direct, with predictable a priori computational complexity.
@@
 This does not mean that the key observation in the minimisation process depends on one particular method for solving the underlying partial differential equations.
-Extension to high-genus surfaces remains a future challenge.
+Extension to closed high-genus surfaces remains a future challenge.
 Automatic detection of charts for an atlas is a very difficult problem and we have not made any attempts to address this question.
@@
 H. Hakula was supported by the Research Council of Finland (Flagship of Advanced Mathematics for Sensing Imaging and Modelling grant 359181).
-A. Rasila and Y. Zheng were supported by NSF of Guangdong Province (No. 2024A1515010467) and Li Ka Shing Foundation GTIIT-STU Joint Research Grant (No. 2024LKSFG06)
+A. Rasila and Y. Zheng were supported by NSF of Guangdong Province (No. 2024A1515010467), Li Ka Shing Foundation GTIIT-STU Joint Research Grant (No. 2024LKSFG06), and GTIIT Education Foundation.